Show your Solution.





1. Two charges, of π‘ž1=+ 3.50 𝑝𝐢 and π‘ž2=βˆ’ 4.50 𝑝𝐢 are 0.400 m apart. Where along the line joining them will net electric field be zero? Note: 𝐸⃗ 𝑛𝑒𝑑= 0 and 𝐸⃗ 1 and 𝐸⃗ 2 are opposite.






2. Four point charges are positioned as such: A(0,1), B(2,1), C(0,0), D(2,0). Charge A = -3 microcoulombs, Charge B = 3 microcoulombs, C = 2 microcoulombs, and D = -2 microcoulombs. What is the net electric field experienced at the origin.

1
Expert's answer
2022-04-11T16:55:39-0400

1.

"x=l\sqrt{\frac{q_2}{q_1}-1}=0.21~m."

2.

"E=E_A+E_B+E_C+E_D,"

"E=k(\frac{q_A}{r_A^2}+\frac{q_B}{r_B^2}+\frac{q_C}{r_C^2}+\frac{q_D}{r_D^2})=3.5~\frac{kN}C."